# Quantifying Data Contamination in Psychometric Evaluations of LLMs

### v1

@@ Abstract
Recent studies apply psychometric questionnaires to Large Language Models (LLMs) to assess high-level psychological constructs such as values, personality, moral foundations, and dark traits. Although prior work has raised concerns about possible data contamination from psychometric inventories, which may threaten the reliability of such evaluations, there has been no systematic attempt to quantify the extent of this contamination. To address this gap, we propose a framework to systematically measure data contamination in psychometric evaluations of LLMs, evaluating three aspects: (1) item memorization, (2) evaluation memorization, and (3) target score matching. Applying this framework to 21 models from major families and four widely used psychometric inventories, we provide evidence that popular inventories such as the Big Five Inventory (BFI-44) and Portrait Values Questionnaire (PVQ-40) exhibit strong contamination, where models not only memorize items but can also adjust their responses to achieve specific target scores. [^1] We will release our code upon publication of the paper.
@@
Jongwook Han, Woojung Song, Jonggeun Lee, Yohan Jo Graduate School of Data Science, Seoul National University {johnhan00, opusdeisong, jonggeun.lee, yohan.jo}@snu.ac.kr
@@ 2.1 Item Memorization
We define item memorization as an LLM’s memorization of items from a psychometric inventory. To evaluate this, we design two tasks: Verbatim Memorization and Key Information Memorization.
Verbatim Memorization.
Verbatim memorization occurs when an LLM reproduces exact text sequences from its training data Carlini et al. (2023). To measure this, we compute the character-level edit distance between model outputs and reference items, following prior work Dong et al. (2024). Formally, for two strings $a$ and $b$, the edit distance $\operatorname{ED}(a,b)$ is defined as the minimum number of single-character insertions, deletions, or substitutions required to transform $a$ into $b$.
We then calculate the normalized Average Edit Distance (AED) across all $N$ items as
$\operatorname{AED}=\frac{1}{N}\sum_{n=1}^{N}\frac{\operatorname{ED}(i_{n},M(p))}{\bar{L}},$
where $i_{n}$ is the $n$-th inventory item, $M(p)$ is the model output given prompt $p$ containing the inventory name and item index $n$, and $\bar{L}$ denotes the average character length of all items. Lower $\operatorname{AED}$ indicates stronger verbatim memorization.
@@
We also evaluate whether LLMs memorize the key semantic content of inventory items. For each item $i_{n}$, we mask the most informative keyword (e.g., in the first item of BFI-10, “I see myself as someone who is reserved”, the word “reserved” is masked). Keywords were annotated by a psychology expert, and the model is prompted with the masked item to generate the missing keyword. The success rate is computed as $\frac{1}{N}\sum_{n=1}^{N}\mathbf{I}(k_{n}=\hat{k}_{n})$, where $N$ is the number of items, $k_{n}$ is the gold keyword, $\hat{k}_{n}$ is the model prediction, and $\mathbf{I}(\cdot)$ is the indicator function. A higher success rate indicates stronger memorization of the core meaning of inventory items.
@@ 3 Results
We evaluate psychometric contamination on four widely used questionnaires: BFI-44, PVQ-40, MFQ, and SD-3, which measure personality traits, human values, moral foundations, and dark triad behaviors. These inventories are selected because they have been frequently adopted in recent NLP studies evaluating LLMs. We evaluate 21 models. The complete list of models, questionnaires, and prompt templates is provided in Appendix B.
@@
Verbatim
Memorization
(AED $\downarrow$)
Key Information
Memorization
(Success Rate $\uparrow$)
Item-Dimension
Mapping
(F1 $\uparrow$)
Option-Score
Mapping
(MAE $\downarrow$)
(MAE $\downarrow$)
@@
gpt-4o-mini 0.82 0.33 0.94 0.72 0.54
gpt-4o 1.06 0.40 0.98 0.09 0.72
gpt-4.1-nano 0.78 0.30 0.92 0.75 1.23
gpt-4.1-mini 0.77 0.35 0.94 0.14 0.62
gpt-4.1 0.72 0.45 0.99 0.28 0.46
gpt-5-nano 2.12 0.33 0.96 0.15 0.12
gpt-5-mini 0.79 0.50 0.98 0.05 0.12
gpt-5 1.34 0.79 1.00 0.02 0.11
@@
qwen3-14b 1.50 0.28 0.94 0.17 0.15
qwen3-32b 1.59 0.29 0.94 0.23 0.20
qwen3-235b-a22b 0.80 0.33 0.95 0.20 0.12
@@
gemini-2.0-flash 0.80 0.46 0.86 0.22 0.47
gemini-2.5-flash-lite 1.97 0.33 0.94 0.32 0.87
gemini-2.5-flash 0.73 0.44 0.96 0.15 0.51
@@
claude-3.5-sonnet 3.98 0.66 0.99 0.05 0.29
claude-4.5-sonnet 8.51 0.64 0.97 0.17 0.45
@@
llama-3.1-70b 2.65 0.35 0.89 1.07 0.65
llama-3.1-405b 6.94 0.42 0.96 1.67 0.76
Average 1.92 0.41 0.95 0.33 0.44
Table 1: Psychometric contamination results averaged across BFI-44, MFQ, PVQ-40, and SD-3. Lower AED and MAE values and a higher success rate and F1 indicate stronger contamination.
@@
The verbatim memorization task yields an average AED of 1.9, indicating that the character-level distance between the generated and original items is nearly twice the length of the original text. This large deviation arises from models paraphrasing items or refusing to output the content due to copyright concerns. In contrast, key information memorization reveals higher semantic retention (success rate $\approx$ 0.4–0.5), suggesting that models are aware of the core semantic content of psychometric items.
@@
Understanding of the evaluation protocol appears to have saturated for most models, with item-dimension mapping achieving an average F1-score of 0.95 across all 21 models. This indicates that LLMs already know which psychological construct each item measures. In option–score mapping, higher-performing models (e.g., GPT-5, GLM-4.5) reach near-zero MAE, showing that they have internalized scoring schemes, including reverse-coded items.
@@
On average, contamination appears to be saturated for recognizing item-dimension mapping and scoring knowledge. Recent model families such as GPT-5, Qwen-3, and GLM exhibit the most consistent contamination pattern, while Claude and Llama variants show higher AED values, largely due to refusal behaviors when reproducing inventory items. As shown in Table 2, PVQ-40 and BFI-44 display stronger contamination compared to MFQ and SD-3, likely reflecting their greater online availability and more frequent use in NLP research.
@@ A.2 Applying Psychometric Inventories in LLM Evaluation
Miotto et al. (2022) applied the HEXACO personality scale and Human Values Scale to GPT-3, showing that the model exhibits personality profiles similar to human respondents. Hadar Shoval et al. (2024) used the Big Five Inventory (BFI), while Tlaie (2024) employed the Moral Foundations Questionnaire (MFQ) to examine moral on LLMs to examine moral and personality-related response patterns. Similarly, Lin et al. (2024) used the Portrait Values Questionnaire (PVQ) to analyze value preferences in LLMs.
@@ B.1 Models
We evaluate models from the following major families: GPT-4 (4o-mini, 4o, 4.1-nano, 4.1-mini, 4.1), GPT-5 (nano, mini, base), Qwen3 (32B, 235B), GLM (4-32B, 4.5 Air, 4.5), Gemini (2.0 Flash, 2.5 Flash Lite, 2.5 Flash), Claude (3.5 Sonnet, 4.5 Sonnet), and Llama 3.1 (70B, 405B). All models are evaluated in their instruction-tuned variants when available, with the temperature fixed to 0. All experiments are conducted via the OpenRouter API, with a total cost of approximately 50 USD.
@@ B.2 Prompts
Listing 1: Prompt : Prompt template for verbatim memorization.
@@
Listing 2: Prompt : Prompt template for keyitem generation.
@@
Listing 3: Prompt : Prompt template for item-dimension mapping.
@@
Listing 4: Prompt : Prompt template for option-score mapping.
@@
Listing 5: Prompt : Prompt template for target score mapping.
@@ Appendix C Additional Results and Analyses
Questionnaire
Verbatim
Memorization
(AED $\downarrow$)
Key Information
Memorization
(Success Rate $\uparrow$)
Item-Dimension
Mapping
(F1 $\uparrow$)
Option-Score
Mapping
(MAE $\downarrow$)
Target Score
Matching
(MAE $\downarrow$)
BFI-44 1.584 0.360 0.978 0.313 0.345
MFQ 2.202 0.375 0.950 0.615 0.697
PVQ-40 1.155 0.587 0.920 0.123 0.354
SD-3 2.726 0.312 0.955 0.255 0.383
@@
Model
Verbatim
Memorization
(AED $\downarrow$)
Key Information
Memorization
(Success Rate $\uparrow$)
Item-Dimension
Mapping
(F1 $\uparrow$)
Option-Score
Mapping
(MAE $\downarrow$)
Target Score
Matching
(MAE $\downarrow$)
@@
gpt-4o-mini 0.71 0.18 0.96 0.27 0.56
gpt-4o 0.77 0.34 1.00 0.00 0.36
gpt-4.1-nano 0.69 0.18 0.93 1.60 0.70
gpt-4.1-mini 0.85 0.25 0.98 0.00 0.65
gpt-4.1 0.76 0.43 1.00 0.75 0.36
gpt-5-nano 2.03 0.20 0.98 0.00 0.00
gpt-5-mini 0.39 0.57 1.00 0.00 0.00
gpt-5 0.83 0.82 1.00 0.07 0.00
@@
qwen3-14b 1.07 0.21 0.98 0.00 0.04
qwen3-32b 0.86 0.19 0.96 0.00 0.00
qwen3-235b-a22b 0.76 0.23 0.98 0.00 0.00
@@
gemini-2.0-flash 0.79 0.52 0.98 0.16 0.32
gemini-2.5-flash-lite 1.29 0.20 0.98 0.87 1.09
gemini-2.5-flash 0.73 0.45 0.98 0.05 0.32
@@
claude-3.5-sonnet 4.89 0.77 1.00 0.00 0.21
claude-4.5-sonnet 9.93 0.66 1.00 0.00 0.43
@@
llama-3.1-70b 1.27 0.18 0.96 1.43 0.85
llama-3.1-405b 2.83 0.34 0.98 1.30 0.48
Table 3: Psychometric contamination results (BFI-44).
Model
Verbatim
Memorization
(AED $\downarrow$)
Key Information
Memorization
(Success Rate $\uparrow$)
Item-Dimension
Mapping
(F1 $\uparrow$)
Option-Score
Mapping
(MAE $\downarrow$)
Target Score
Matching
(MAE $\downarrow$)

### v2

@@ Abstract
Recent studies apply psychometric questionnaires to Large Language Models (LLMs) to assess high-level psychological constructs such as values, personality, moral foundations, and dark traits. Although prior work has raised concerns about possible data contamination from psychometric inventories, which may threaten the reliability of such evaluations, there has been no systematic attempt to quantify the extent of this contamination. To address this gap, we propose a framework to systematically measure data contamination in psychometric evaluations of LLMs, evaluating three aspects: (1) item memorization, (2) evaluation memorization, and (3) target score matching. Applying this framework to 21 models from major families and four widely used psychometric inventories, we provide evidence that popular inventories such as the Big Five Inventory (BFI-44) and Portrait Values Questionnaire (PVQ-40) exhibit strong contamination, where models not only memorize items but can also adjust their responses to achieve specific target scores. [^1] Our code is available at https://github.com/holi-lab/psychometric-contamination.
@@
Jongwook Han    Woojung Song    Jonggeun Lee    Yohan Jo Graduate School of Data Science, Seoul National University {johnhan00,opusdeisong,jonggeun.lee,yohan.jo}@snu.ac.kr
@@ 2.1 Item Memorization
We define item memorization as an LLM’s memorization of items from a psychometric inventory. To evaluate this, we design two tasks: Semantic Memorization and Key Information Memorization.
Semantic Memorization.
To analyze whether LLMs have prior information about psychometric items, we evaluate their ability to reconstruct item content when provided with the inventory name and item index. We prompt models to generate the corresponding item text. Then we filter out refusal responses using GPT-5.2-mini as an automated filter. For the remaining responses, we compute the embedding similarity between the original and the generated item text. We use OpenAI’s text-embedding-3-large model and compute cosine similarity. We focus on semantic similarity rather than exact text overlap, as models may paraphrase item content.
@@
We also evaluate whether LLMs memorize the key semantic content of inventory items. For each item $i_{n}$, we mask the most informative keyword (e.g., in the first item of BFI-10, “I see myself as someone who is reserved”, the word “reserved” is masked). Keywords were annotated by a psychology expert, and the model is prompted with the masked item to generate the missing keyword. To validate the the reliability of the keyword annotations, we collected three additional human annotations and computed the Fleiss’ Kappa among the four annotators. The resulting agreement scores were 0.9187 for BFI-44, 0.7893 for MFQ, 0.8168 for PVQ-40, and 0.7548 for SD3. These results indicate that the inter-annotator agreement is consistently high across inventories, suggesting that the masked-keyword annotations are reliable.
We compute the success rate as $\frac{1}{N}\sum_{n=1}^{N}\mathbf{I}(k_{n}=\hat{k}_{n})$, where $N$ is the number of items, $k_{n}$ is the gold keyword, $\hat{k}_{n}$ is the model prediction, and $\mathbf{I}(\cdot)$ is the indicator function. A higher success rate indicates stronger memorization of the core meaning of inventory items.
@@ 3 Results
We evaluate psychometric contamination on four widely used questionnaires: BFI-44, PVQ-40, MFQ, and SD-3, which measure personality traits, human values, moral foundations, and dark triad behaviors. These inventories are selected because they have been frequently adopted in recent NLP studies evaluating LLMs. We evaluate 21 models. The complete list of models, questionnaires, and prompt templates is provided in Appendix B. For all experiments we set the temperature to 0.7, repeat three times, and report the averaged results alongside 95 percent confidence intervals.
@@
Semantic Memorization (Similarity $\uparrow$) Key Information Memorization (Success Rate $\uparrow$) Item-Dimension Mapping (F1 $\uparrow$) Option-Score Mapping (MAE $\downarrow$) (MAE $\downarrow$)
@@
gpt-4o-mini 0.23 $\pm$ 0.01 0.33 $\pm$ 0.04 0.95 $\pm$ 0.03 0.73 $\pm$ 0.24 0.53 $\pm$ 0.04
gpt-4o 0.40 $\pm$ 0.02 0.40 $\pm$ 0.06 0.98 $\pm$ 0.01 0.10 $\pm$ 0.13 0.77 $\pm$ 0.17
gpt-4.1-nano 0.21 $\pm$ 0.01 0.27 $\pm$ 0.07 0.92 $\pm$ 0.07 0.85 $\pm$ 0.29 1.21 $\pm$ 0.08
gpt-4.1-mini 0.30 $\pm$ 0.01 0.33 $\pm$ 0.11 0.93 $\pm$ 0.02 0.21 $\pm$ 0.28 0.69 $\pm$ 0.11
gpt-4.1 0.36 $\pm$ 0.02 0.46 $\pm$ 0.03 0.99 $\pm$ 0.00 0.32 $\pm$ 0.13 0.44 $\pm$ 0.12
gpt-5-nano 0.28 $\pm$ 0.05 0.34 $\pm$ 0.12 0.96 $\pm$ 0.03 0.12 $\pm$ 0.07 0.10 $\pm$ 0.06
gpt-5-mini 0.43 $\pm$ 0.02 0.54 $\pm$ 0.03 0.98 $\pm$ 0.03 0.05 $\pm$ 0.00 0.12 $\pm$ 0.03
gpt-5 0.56 $\pm$ 0.02 0.79 $\pm$ 0.10 1.00 $\pm$ 0.01 0.01 $\pm$ 0.04 0.10 $\pm$ 0.02
@@
qwen3-14b 0.20 $\pm$ 0.01 0.29 $\pm$ 0.09 0.94 $\pm$ 0.03 0.21 $\pm$ 0.11 0.15 $\pm$ 0.11
qwen3-32b 0.21 $\pm$ 0.01 0.30 $\pm$ 0.10 0.96 $\pm$ 0.05 0.28 $\pm$ 0.25 0.23 $\pm$ 0.12
qwen3-235b-a22b 0.34 $\pm$ 0.01 0.18 $\pm$ 0.03 0.78 $\pm$ 0.19 1.06 $\pm$ 0.69 0.12 $\pm$ 0.06
@@
gemini-2.0-flash-001 0.32 $\pm$ 0.01 0.45 $\pm$ 0.08 0.82 $\pm$ 0.04 0.21 $\pm$ 0.05 0.44 $\pm$ 0.10
gemini-2.5-flash-lite 0.25 $\pm$ 0.01 0.33 $\pm$ 0.08 0.93 $\pm$ 0.03 0.32 $\pm$ 0.00 0.81 $\pm$ 0.07
gemini-2.5-flash 0.33 $\pm$ 0.01 0.41 $\pm$ 0.09 0.95 $\pm$ 0.05 0.18 $\pm$ 0.24 0.56 $\pm$ 0.10
@@
claude-3.5-sonnet 0.34 $\pm$ 0.13 0.65 $\pm$ 0.10 0.99 $\pm$ 0.00 0.05 $\pm$ 0.03 0.35 $\pm$ 0.05
claude-sonnet-4.5 0.46 $\pm$ 0.02 0.64 $\pm$ 0.03 0.98 $\pm$ 0.01 0.16 $\pm$ 0.01 0.47 $\pm$ 0.02
@@
llama-3.1-70b-instruct 0.10 $\pm$ 0.02 0.35 $\pm$ 0.11 0.91 $\pm$ 0.05 1.34 $\pm$ 0.54 0.69 $\pm$ 0.13
llama-3.1-405b-instruct 0.16 $\pm$ 0.03 0.33 $\pm$ 0.10 0.88 $\pm$ 0.07 2.29 $\pm$ 0.45 0.60 $\pm$ 0.15
Average 0.31 0.39 0.94 0.44 0.45
Table 1: Psychometric contamination results averaged across BFI-44, MFQ, PVQ-40, and SD-3. Values are reported as mean $\pm$ 95% confidence interval (CI) across three evaluation runs. Higher cosine similarity, success rate, and F1, and lower MAE indicate stronger contamination.
@@
The semantic memorization task yields an average embedding similarity of 0.31. The key information memorization reveals higher semantic retention (average success rate of 0.39), suggesting that models are aware of the core semantic content of psychometric items.
@@
Understanding of the evaluation protocol appears to have saturated for most models, with item-dimension mapping achieving an average F1-score of 0.94 across all 21 models. This indicates that LLMs already know which psychological construct each item measures. In option–score mapping, higher-performing models (e.g., GPT-5, GLM-4.5) exhibit low MAE, showing that they have internalized scoring schemes, including reverse-coded items.
@@
On average, contamination appears to be saturated for recognizing item-dimension mapping and scoring knowledge. Recent model families such as GPT-5, Claude, and GLM exhibit the most consistent contamination pattern, while Llama variants show less. As shown in Table 2, PVQ-40 and BFI-44 display stronger contamination compared to MFQ and SD-3, likely reflecting their greater online availability and more frequent use in NLP research. For more detailed analysis, see Appendix C.
@@
Recommendations.
Because LLMs may have been exposed to many psychometric inventories available online, directly comparing the magnitude of contamination signals across settings is not straightforward. To facilitate interpretation, we define simple baselines for each contamination setting and provide practical guidelines for interpreting the results.
For semantic memorization, we use dimension description text as a baseline (i.e., a short definition of the trait/value that the item is measuring) and measure its embedding similarity to the item text. This baseline yields cosine similarities of 0.32 (BFI-44), 0.30 (MFQ), 0.41 (PVQ-40), and 0.19 (SD3), with an overall average of 0.32 across inventories. If a model’s semantic memorization score is close to the baseline, it suggests that the model’s reconstruction is at the level of generic dimension definition.
For key information memorization, establishing a baseline corresponding to a truly uncontaminated model is inherently challenging, since there is no well-defined procedure for approximating random keyword generation under conditions of potential contamination. We therefore contextualize the results in Table 1 using the empirical distribution of performance across the 21 evaluated models. Specifically, under the assumption that lower-scoring models are likely to exhibit less contamination, we report the 10th-percentile success rates: 0.16 (BFI-44), 0.25 (MFQ), 0.41 (PVQ-40), and 0.18 (SD-3). These values should be interpreted cautiously, however, as even low-scoring models may exhibit some degree of contamination. We adopt the 10th percentile to reduce sensitivity to outliers, but this choice is inherently heuristic, and alternative percentile thresholds could be adopted depending on the practitioner’s objectives.
Regarding absolute values, the success rate has a direct operational meaning where 0 indicates that the model never recovers the important keyword of an item, whereas 1 indicates that it recovers it for all items. Intermediate values can be interpreted as frequency. For example, a score of 0.5 means that the model knows the exact wording of the important keyword for about half of the items, suggesting that the model knows the semantics of at least half of the items in an inventory.
For item-dimension mapping, a random classifier over $N$ dimensions yields an expected F1 of $1/N$ (0.2 for BFI-44 with $N=5$). An F1 score near the baseline suggests the model is simply guessing.
For option-score mapping, if a model assigns scores uniformly at random on a $K$-point scale inventory, the expected mean absolute error is $(K^{2}-1)/(3K)$ (1.60 for $K=5$ and 1.94 for $K=6$). MAE score near this baseline indicates the model is guessing the score associated with each response option.
For target score matching, we use a random baseline where the model selects a response uniformly at random from a $K$-point scale inventory, independent of the target. Under this strategy, the expected MAE is $\mathbb{E}[|t-\hat{t}|]=\frac{(t-1)t+(K-t)(K-t+1)}{2K}$, where $t$ is the target score and $\hat{t}$ is the randomly selected score. Averaging over the three target conditions (minimum, mean, and maximum) yields expected MAE values of 1.73 for $K=5$ and 2.17 for $K=6$. An MAE close to the baseline indicates that the model is guessing randomly.
@@ 4 Conclusion
5 Acknowledgments
This work was supported by the Creative-Pioneering Researchers Program through Seoul National University and by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2024-00333484). It was also supported by the National Supercomputing Center with supercomputing resources including technical support (KSC-2025-CRE-0332, KSC-2025-CRE-0514).
@@ A.2 Applying Psychometric Inventories in LLM Evaluation
Miotto et al. (2022) applied the HEXACO personality scale and Human Values Scale to GPT-3, showing that the model exhibits personality profiles similar to human respondents. Hadar Shoval et al. (2024) used the Big Five Inventory (BFI), while Tlaie (2024) employed the Moral Foundations Questionnaire (MFQ) to examine moral and personality-related response patterns in LLMs. Similarly, Lin et al. (2024) used the Portrait Values Questionnaire (PVQ) to analyze value preferences in LLMs.
@@ B.1 Models
We evaluate models from the following major families: GPT-4 (4o-mini, 4o, 4.1-nano, 4.1-mini, 4.1), GPT-5 (nano, mini, base), Qwen3 (32B, 235B), GLM (4-32B, 4.5 Air, 4.5), Gemini (2.0 Flash, 2.5 Flash Lite, 2.5 Flash), Claude (3.5 Sonnet, 4.5 Sonnet), and Llama 3.1 (70B, 405B). All models are evaluated in their instruction-tuned variants when available, with the temperature set to 0.7. All experiments are conducted via the OpenRouter API, with a total cost of approximately 50 USD.
@@ B.2 Prompts
Listing 1: Prompt 1: Prompt template for semantic memorization.
You are given two texts: (1) the original text and (2) the candidate text.
Your task is to evaluate the candidate text according to two criteria.
Task 1 -- Refusal Detection
Determine whether the candidate text is a refusal.
A refusal means the model declines, deflects, or avoids performing the task (e.g., I cannot answer, Im unable to help, I dont have enough information).
Respond with Yes or No.
Task 2 -- Similarity Evaluation
Evaluate how similar the candidate text is to the original text.
Use the following options exactly as written:
Highly Similar, Moderately Similar, Slightly Similar, Slightly Dissimilar, Moderately Dissimilar, Highly Dissimilar
Similarity should consider meaning, intent, and content---not surface-level phrasing.
Original text: {original}
Candidate text: {candidate}
Listing 2: Prompt 2: Prompt template for refusal detection and similarity evaluation.
@@
Listing 3: Prompt 3: Prompt template for keyitem generation.
@@
Listing 4: Prompt 4: Prompt template for item-dimension mapping.
@@
Listing 5: Prompt 5: Prompt template for option-score mapping.
@@
Listing 6: Prompt 6: Prompt template for target score mapping.
@@ Appendix C Additional Results and Analyses
As mentioned in Section 2.1, we filtered out refusal responses with GPT-5.2-mini. We used Prompt 2 to detect refusal responses. In the same prompt, we also asked the model to grade the similarity between the reference text and the generated text on a 6-point scale, ranging from highly dissimilar to highly similar. Tables 7, 8, 9, and 10 show the refusal rates and similarity scores on the BFI-44, MFQ, PVQ-40, and SD-3 inventories.
Contamination differences across questionnaires.
In Section 3, we observed that PVQ-40 and BFI-44 exhibit stronger contamination signals than MFQ and SD-3, and hypothesized that this difference may be related to their greater online and academic exposure. To further justify this interpretation, we examined the citation counts of the primary papers associated with each inventory. The results are as follows: BFI (12,560) Rammstedt and John (2007), PVQ (7,087) Schwartz (2012), MFQ (4,226) Graham et al. (2011), and SD3 (3,369) Jones and Paulhus (2013), as of 2025. While citation counts are an approximate indicator of exposure, these differences suggest that BFI and PVQ have broader academic visibility than MFQ and SD-3, which may contribute to greater exposure during model training.
Baseline settings.
For the semantic memorization baseline, we use the dimension description text for each inventory. For BFI-44, we use the trait descriptions from Pervin and John (1999). For MFQ, we use the foundation descriptions provided on https://moralfoundations.org/. For PVQ-40, we use the value descriptions from Schwartz (2012). For SD-3, we use dimension descriptions adapted from Jones and Paulhus (2013).
#BFI-44
Openness: Ideas (curious), fantasy (imaginative), aesthetics (artistic), actions (wide interests), feelings (excitable), and values (unconventional).
Conscientiousness: Competence (efficient), order (organized), dutifulness (not careless), achievement striving (thorough), self-discipline (not lazy), and deliberation (not impulsive).
Extraversion: Gregariousness (sociable), assertiveness (forceful), activity (energetic), excitement-seeking (adventurous), positive emotions (enthusiastic), and warmth (outgoing).
Agreeableness: Trust (forgiving), straightforwardness (not demanding), altruism (warm), compliance (not stubborn), modesty (not show-off), and tender-mindedness (sympathetic).
Neuroticism: Anxiety (tense), angry hostility (irritable), depression (not contented), self-consciousness (shy), impulsiveness (moody), and vulnerability (not self-confident).
#MFQ
Care: This foundation is related to our long evolution as mammals with attachment systems and an ability to feel (and dislike) the pain of others. It underlies the virtues of kindness, gentleness, and nurturance.
Fairness: This foundation is related to the evolutionary process of reciprocal altruism. It underlies the virtues of justice and rights.
Loyalty: This foundation is related to our long history as tribal creatures able to form shifting coalitions. It is active anytime people feel that its one for all and all for one. It underlies the virtues of patriotism and self-sacrifice for the group.
Authority: This foundation was shaped by our long primate history of hierarchical social interactions. It underlies virtues of leadership and followership, including deference to prestigious authority figures and respect for traditions.
Purity: This foundation was shaped by the psychology of disgust and contamination. It underlies notions of striving to live in an elevated, less carnal, more noble, and more natural way (often present in religious narratives). This foundation underlies the widespread idea that the body is a temple that can be desecrated by immoral activities and contaminants (an idea not unique to religious traditions). It underlies the virtues of self-discipline, self-improvement, naturalness, and spirituality.
#PVQ-40
Self-Direction: Independent thought and action--choosing, creating, and exploring.
Stimulation: Excitement, novelty, and challenge in life.
Hedonism: Pleasure or sensuous gratification for oneself.
Achievement: Personal success through demonstrating competence according to social standards.
Power: Social status and prestige, control or dominance over people and resources.
Security: Safety, harmony, and stability of society, of relationships, and of self.
Conformity: Restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms.
Tradition: Respect, commitment, and acceptance of the customs and ideas that ones culture or religion provides.
Benevolence: Preserving and enhancing the welfare of those with whom one is in frequent personal contact (the in-group’).
Universalism: Understanding, appreciation, tolerance, and protection for the welfare of all people and for nature.
#SD-3
Machiavellianism: Manipulativeness, callous affect, and a strategic-calculating orientation.
Psychopathy: Deficits in affect (i.e., callousness) and self-control (i.e., impulsivity).
Narcissism: Ego-identity goals drive narcissistic behavior.
Listing 7: Dimension definitions of inventories used in semantic memorization baseline experiments.
Questionnaire Semantic Memorization (Similarity $\uparrow$) Key Information Memorization (Success Rate $\uparrow$) Item-Dimension Mapping (F1 $\uparrow$) Option-Score Mapping (MAE $\downarrow$) Target Score Matching (MAE $\downarrow$)
BFI-44 0.36 0.34 0.96 0.41 0.37
MFQ 0.17 0.35 0.95 0.70 0.68
PVQ-40 0.37 0.57 0.91 0.28 0.35
SD-3 0.24 0.31 0.93 0.37 0.41
@@
Semantic Memorization (Similarity $\uparrow$) Key Information Memorization (Success Rate $\uparrow$) Item–Dimension Mapping (F1 $\uparrow$) Option–Score Mapping (MAE $\downarrow$) (MAE $\downarrow$)
@@
gpt-4o-mini 0.35 $\pm$ 0.02 0.17 $\pm$ 0.09 0.96 $\pm$ 0.00 0.25 $\pm$ 0.08 0.56 $\pm$ 0.04
gpt-4o 0.44 $\pm$ 0.04 0.36 $\pm$ 0.06 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.38 $\pm$ 0.11
gpt-4.1-nano 0.37 $\pm$ 0.02 0.16 $\pm$ 0.06 0.92 $\pm$ 0.06 1.61 $\pm$ 0.29 0.73 $\pm$ 0.12
gpt-4.1-mini 0.21 $\pm$ 0.02 0.20 $\pm$ 0.15 0.98 $\pm$ 0.00 0.02 $\pm$ 0.10 0.69 $\pm$ 0.18
gpt-4.1 0.25 $\pm$ 0.03 0.41 $\pm$ 0.00 1.00 $\pm$ 0.00 0.83 $\pm$ 0.21 0.36 $\pm$ 0.14
gpt-5-nano 0.40 $\pm$ 0.07 0.25 $\pm$ 0.11 0.98 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.01
gpt-5-mini 0.58 $\pm$ 0.04 0.57 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
gpt-5 0.67 $\pm$ 0.02 0.84 $\pm$ 0.11 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
@@
qwen3-14b 0.28 $\pm$ 0.02 0.21 $\pm$ 0.03 0.98 $\pm$ 0.00 0.00 $\pm$ 0.00 0.03 $\pm$ 0.08
qwen3-32b 0.29 $\pm$ 0.01 0.18 $\pm$ 0.11 0.97 $\pm$ 0.04 0.00 $\pm$ 0.00 0.04 $\pm$ 0.09
qwen3-235b-a22b 0.28 $\pm$ 0.03 0.14 $\pm$ 0.06 0.81 $\pm$ 0.16 0.86 $\pm$ 0.59 0.00 $\pm$ 0.00
@@
gemini-2.0-flash-001 0.27 $\pm$ 0.02 0.52 $\pm$ 0.06 0.98 $\pm$ 0.00 0.16 $\pm$ 0.14 0.41 $\pm$ 0.07
gemini-2.5-flash-lite 0.26 $\pm$ 0.02 0.21 $\pm$ 0.03 0.98 $\pm$ 0.00 0.87 $\pm$ 0.00 1.04 $\pm$ 0.08
gemini-2.5-flash 0.22 $\pm$ 0.02 0.39 $\pm$ 0.11 0.98 $\pm$ 0.06 0.05 $\pm$ 0.00 0.36 $\pm$ 0.07
@@
claude-3.5-sonnet 0.29 $\pm$ 0.37 0.73 $\pm$ 0.11 1.00 $\pm$ 0.02 0.00 $\pm$ 0.00 0.28 $\pm$ 0.08
claude-sonnet-4.5 0.53 $\pm$ 0.09 0.67 $\pm$ 0.03 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.46 $\pm$ 0.04
@@
llama-3.1-70b-instruct 0.21 $\pm$ 0.08 0.19 $\pm$ 0.03 0.96 $\pm$ 0.03 1.40 $\pm$ 0.79 0.87 $\pm$ 0.22
llama-3.1-405b-instruct 0.20 $\pm$ 0.30 0.23 $\pm$ 0.03 0.87 $\pm$ 0.07 2.25 $\pm$ 0.17 0.53 $\pm$ 0.06
Average 0.36 0.34 0.96 0.41 0.37
Table 3: Psychometric contamination results for BFI-44 with 95% confidence intervals from three experimental repetitions.
Semantic Memorization (Similarity $\uparrow$) Key Information Memorization (Success Rate $\uparrow$) Item–Dimension Mapping (F1 $\uparrow$) Option–Score Mapping (MAE $\downarrow$) (MAE $\downarrow$)